### v2

@@ Abstract
Current research on trajectory prediction primarily relies on data collected by onboard sensors of an ego vehicle. With the rapid advancement in connected technologies, such as vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication, valuable information from alternate views becomes accessible via wireless networks. The integration of information from alternative views has the potential to overcome the inherent limitations associated with a single viewpoint, such as occlusions and limited field of view. In this work, we introduce V2INet, a novel trajectory prediction framework designed to model multi-view data by extending existing single-view models. Unlike previous approaches where the multi-view data is manually fused or formulated as a separate training stage, our model supports end-to-end training, enhancing both flexibility and performance. Moreover, the predicted multimodal trajectories are calibrated by a post-hoc conformal prediction module to get valid and efficient confidence regions. We evaluated the entire framework using the real-world V2I dataset V2X-Seq. Our results demonstrate superior performance in terms of Final Displacement Error (FDE) and Miss Rate (MR) using a single GPU. The code is publicly available at: https://github.com/xichennn/V2I_trajectory_prediction.
@@ 1 Introduction
Trajectory prediction plays a critical role in autonomous driving. Typically, we rely on the on-board sensors of an ego vehicle to gather environmental information necessary for performing various autonomous driving tasks. However, with the rapid advancement in connected technologies, such as vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication, valuable information from alternate views becomes accessible via wireless networks. The integration of information from alternative views has the potential to overcome the inherent limitations associated with a single viewpoint, such as occlusions and limited field of view. We categorize the data obtained solely from ego vehicle on-board sensors or infrastructure as single-view information, whereas the accessibility to multiple viewpoints is referred to as multi-view data. Figure 1 depicts an intersection scenario where the AV faces a potential left-turn collision due to its field of view being obstructed by a large truck. Roadside cameras, strategically positioned to have an unobstructed view of the entire intersection, provide a comprehensive overview of the traffic situation. This enhanced perspective allows the AV to receive critical information and effectively avoid the impending collision.
@@
From a trajectory predictive modeling perspective, significant research efforts have been dedicated to leveraging single-view datasets collected by the ego vehicle [1, 2, 3, 4, 5]. These efforts typically involve modeling the temporal dependencies, agent-agent interactions and agent-lane relations. However, the challenge arises when dealing with multi-view data, particularly in how to effectively fuse the information due to overlapping field of views.
@@
We introduce a trajectory prediction framework that utilizes multi-view data without the need for explicit association between different perspectives. Rather than developing a specialized multi-view model, our approach seamlessly integrates with state-of-the-art single-view trajectory models, maximizing the utility of existing research efforts. No special training strategies are required. We can easily leverage the pretrained single-view trajectory models to expediate the training. Specifically, for trajectory data from each single-view, we employ established graph neural network (GNN) based models, such as LaneGCN [1], HiVT [3], to capture temporal dependencies, agent-agent interactions, and agent-lane relations. Then we utilize a cross-graph attention module to fuse the node embeddings from different views. The fused final embeddings will then go through a multimodal decoder to get future trajectory predictions.
Existing works have modeled the multi-modality explicitly by introducing anchors [2, 8, 9], mixture models [10, 11, 12, 13, 1, 3], or implicitly through latent variables such as Conditional Variational Auto-encoder (CVAE) or generative models [14, 15, 16]. The implicit models often face the issue of mode collapse, therefore we will model the multi-modality by a mixture model built upon MLP. It is a common strategy to assign a higher score to the modality closer to the ground truth during training. This strategy, however, may encounter robustness issues during inference. Figure 2 presents two examples where the top-scored prediction is not the closest to the ground truth trajectory.
@@
This arises from the inherent randomness in the ground truth data, posing challenges for the model to learn the priorities of multiple predictions. Relying on the scores to rank the predictions can lead to hazardous results in safety-critical scenarios. Additionally, while most mixture models like GMM provide both point estimates and variance estimation, the reliability of these uncertainty quantification (UQ) remains uncertain. Unfortunately, this aspect is often overlooked in the model evaluation process. However, UQ can offer valuable insights for downstream decision-making processes. To address these challenges, we introduce conformal prediction (CP) [17, 18], as a post-hoc module to our framework. CP creates statistically rigorous uncertainty intervals for the predictions that are guaranteed to contain the ground truth with a user-specified probability.
@@ 2.1 Cooperative Driving Dataset
Collaborative perception has been the most extensively studied area in collaborative driving [19, 20, 21, 22]. It leverages the wireless communication technologies, such as V2V and V2I, to share information, which enables perception beyond line-of-sight and field-of-view hence overcoming common perceptual shortcomings with individual perception, such as blind spot, occlusions and long-range issues [23, 24]. Many datasets have been collected in both simulation environment [25, 26, 27] and real-world [28, 29] to facilitate the study. A comprehensive comparison can be found in Table 1. While majority of them focused on the upstream tasks such as detection, tracking and segmentation, to enable well-informed decision-making for autonomous vehicles, however, it’s critical to also incorporate V2X data for predicting the behavior of surrounding traffic participants. V2X-Seq/Forecasting [6] and V2X-Traj [7] present two real-world datasets specifically designed for cooperative trajectory prediction tasks. The experiments in V2X-Seq/Forecasting demonstrate that leveraging the infrastructure-side trajectories can enhance the trajectory prediction performance. V2X-Traj aims for more general V2X scenarios, extending to V2V cooperation.
@@ 2.3 Uncertainty Quantification
In the trajectory prediction task, numerous sources of uncertainty exist, including inherent multi-modality, partial observability, short time scales, data limitations, intention type imbalances, and domain gaps. Most existing trajectory prediction models address uncertainty by maximizing the likelihood of an assumed distribution, such as Gaussian or Laplace [9, 3, 1, 11]. However, trajectories with the largest likelihood are often nonsensical [37, 38]. Alternatively, some research focuses on approximating Bayesian inference for deep learning models using techniques like Monte Carlo dropout [gal2016dropout], which involves performing stochastic forward passes through the network and averaging the results. Among these approaches, [39] stands out as the only work that incorporates collaborative uncertainty among agents into the modeling process to guide the ranking of multimodal trajectories by uncertainty, albeit requiring special training strategies. However, none of the predicted uncertainties from these methods offer finite sample coverage guarantees, which is suboptimal for safety-critical applications such as vehicle trajectory prediction.
@@ 4.3 Single View Encoder
We exemplify here with HiVT [3] where only attention mechanism is employed, with the adoption of a rotation-invariant representation. To attend to the local information, at each timestamp, the surrounding actors’ information is aggregated and the embeddings at each time stamp subsequently go through a transformer to capture the temporal dependencies. After obtaining the spatio-temporal embeddings, the surrounding lane information is aggregated for each actor. The update actor node embeddings is then passed to a global interaction module which aggregates the local context of different actors and updates each actor’s representation to capture long-range dependencies and scene-level dynamics. Here’s the proofread version:
We denote the final embedding for actor node $i$ from the vehicle view as $\mathbf{h}_{i}$, and actor node $j$ from the infrastructure view as $\mathbf{h}_{j}$. Both $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are in $\mathbb{R}^{d_{h}}$, where $d_{h}$ is the embedding dimension.

### v3

@@ Abstract
Current research on trajectory prediction primarily relies on data collected by onboard sensors of an ego vehicle. With the rapid advancement in connected technologies, such as vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication, valuable information from alternate views becomes accessible via wireless networks. The integration of information from alternative views has the potential to overcome the inherent limitations associated with a single viewpoint, such as occlusions and limited field of view. In this work, we introduce V2INet, a novel trajectory prediction framework designed to model multi-view data by extending existing single-view models. Unlike previous approaches where the multi-view data is manually fused or formulated as a separate training stage, our model supports end-to-end training, enhancing both flexibility and performance. Moreover, the predicted multimodal trajectories are calibrated by a post hoc conformal prediction module to get valid and efficient confidence regions. We evaluated the entire framework on the real-world V2I dataset V2X-Seq. Our results demonstrate superior performance in terms of Final Displacement Error (FDE) and Miss Rate (MR) using a single GPU. The code is publicly available at: https://github.com/xichennn/V2I_trajectory_prediction.
@@ 1 Introduction
Trajectory prediction plays a critical role in autonomous driving. Typically, we rely on the on-board sensors of an ego vehicle to gather surrounding information necessary for performing various autonomous driving tasks. However, with the rapid advancement in connected technologies, such as vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication, valuable information from alternate views becomes accessible via wireless networks. The integration of information from alternative views has the potential to overcome the inherent limitations associated with a single viewpoint, such as occlusions and limited field of view. We categorize the data obtained solely from ego vehicle on-board sensors or infrastructure as single-view information, whereas the accessibility to multiple viewpoints is referred to as multi-view data. Figure 1 depicts an intersection scenario where the AV faces a potential left-turn collision due to its field of view being obstructed by a large truck. Roadside cameras, strategically positioned to have an unobstructed view of the entire intersection, provide a comprehensive overview of the traffic situation. This enhanced perspective allows the AV to receive critical information and effectively avoid the impending collision.
@@
From a trajectory predictive modeling perspective, significant research efforts have been dedicated to only to use single-view datasets collected by the ego vehicle [1, 2, 3, 4, 5]. These efforts typically involve modeling the temporal dependencies, agent-agent interactions, and agent-lane relations. However, the challenge arises when dealing with multi-view data, particularly in how to effectively fuse the information due to overlapping field of views.
@@
We introduce a trajectory prediction framework that utilizes multi-view data without the need for explicit association between different perspectives. Rather than developing a specialized multi-view model, our approach seamlessly integrates with state-of-the-art single-view trajectory models, maximizing the utility of existing research efforts. No special training strategies are required. We can easily take advantage of the pretrained single-view trajectory models to expedite the training. Specifically, for trajectory data from each single-view, we employ established graph neural network (GNN) based models, such as LaneGCN [1], HiVT [3], to capture temporal dependencies, agent-agent interactions, and agent-lane relations. Then we utilize a cross-graph attention module to fuse the node embeddings from different views. The fused final embeddings will then go through a multimodal decoder to get future trajectory predictions.
Existing works have modeled the multi-modality explicitly by introducing anchors [2, 8, 9], mixture models [10, 11, 12, 13, 1, 3], or implicitly through latent variables such as Conditional Variational Auto-encoder (CVAE) or generative models [14, 15, 16]. The implicit models often face the issue of mode collapse, therefore we will model the multi-modality by a mixture model built upon MLP. It is a common strategy to assign a higher score to the modality closer to the ground truth during training. The strategy, however, may encounter robustness issues during inference. Figure 2 presents two examples where the top-scored prediction is not the closest to the ground truth trajectory.
@@
The situation arises from the inherent randomness in the ground truth data, posing challenges for the model to learn the priorities of multiple predictions. Relying on the scores to rank the predictions can lead to hazardous results in safety-critical scenarios. Additionally, while most mixture models like the Gaussian Mixture Model (GMM) provide both point estimation and variance estimation, the reliability of these uncertainty quantifications (UQ) remains questionable. Unfortunately, this aspect is often overlooked in the model evaluation process. However, UQ can offer valuable insights for downstream decision-making processes. To address these challenges, we introduce conformal prediction (CP) [17, 18], as a post-hoc module to our framework. CP creates statistically rigorous uncertainty intervals for the predictions that are guaranteed to contain the ground truth with a user-specified probability.
@@ 2.1 Cooperative Driving Dataset
Collaborative perception has been the most extensively studied area in collaborative driving [19, 20, 21, 22]. It uses the recent developments in wireless communication technologies, such as V2V and V2I, to share information, which enables perception beyond line-of-sight and field-of-view hence overcoming common perceptual shortcomings with individual perception, such as blindspot, occlusions, and long-range issues [23, 24]. Many datasets have been collected in both simulation environment [25, 26, 27] and real-world [28, 29] to facilitate the study. A comprehensive comparison can be found in Table 1. While a majority of them focused on the upstream tasks such as detection, tracking, and segmentation, to enable well-informed decision-making for autonomous vehicles, however, it’s critical to also incorporate V2X data for predicting the behavior of surrounding traffic participants. V2X-Seq/Forecasting [6] and V2X-Traj [7] present two real-world datasets specifically designed for cooperative trajectory prediction tasks. The experiments in V2X-Seq/Forecasting demonstrate that leveraging the infrastructure-side trajectories can enhance the trajectory prediction performance. V2X-Traj aims for more general V2X scenarios, extending to V2V cooperation.
@@ 2.3 Uncertainty Quantification
In the trajectory prediction task, numerous sources of uncertainty exist, including inherent multi-modality, partial observability, short time scales, data limitations, intention type imbalances, and domain gaps. Most existing trajectory prediction models address uncertainty by maximizing the likelihood of an assumed distribution, such as Gaussian or Laplace [9, 3, 1, 11]. However, trajectories with the largest likelihood are often nonsensical [37, 38]. Alternatively, some research focuses on approximating Bayesian inference for deep learning models using techniques like Monte Carlo dropout [39], which involves performing stochastic forward passes through the network and averaging the results. Among these approaches, [40] stands out as the only work that incorporates collaborative uncertainty among agents into the modeling process to guide the ranking of multimodal trajectories by uncertainty, albeit requiring special training strategies. However, none of the predicted uncertainties from these methods offer finite sample coverage guarantees, which is suboptimal for safety-critical applications such as vehicle trajectory prediction.
@@ 4.3 Single View Encoder
We exemplify here with HiVT [3] where only attention mechanism is employed, with the adoption of a rotation-invariant representation. To attend to the local information, at each timestamp, the surrounding actors’ information is aggregated and the embeddings at each time stamp subsequently go through a transformer to capture the temporal dependencies. After obtaining the spatio-temporal embeddings, the surrounding lane information is aggregated for each actor. The update actor node embeddings is then passed to a global interaction module which aggregates the local context of different actors and updates each actor’s representation to capture long-range dependencies and scene-level dynamics. We denote the final embedding for actor node $i$ from the vehicle view as $\mathbf{h}_{i}$, and actor node $j$ from the infrastructure view as $\mathbf{h}_{j}$. Both $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are in $\mathbb{R}^{d_{h}}$, where $d_{h}$ is the embedding dimension.